# Autotune: fast, accurate, and automatic tuning parameter selection for LASSO

### v1

@@ 3.3 An illustration of PR properties
with $n=80$ observations, $p=750$ covariates, and $\beta_{0}=(1,1,1,1,1,0,0,\dots)\in\real^{p}$, i.e., yi’s depend on only the first 5 covariates. The $\text{X}_{i,\cdot}$’s are independently generated from $p$-dimensional standard normal distribution. Fixing $\text{X}_{i,\cdot}$’s, we generate yi’s multiple times by drawing errors $\epsilon_{i}$’s from $\mathcal{N}(0,\sigma^{2})$ with $\sigma^{2}$ chosen so that Signal-to-Noise Ratio is 1.
We exhibit the histogram of the standard deviations (SD) of all the partial residuals ˜4 obtained from the estimates of $\beta$ at the end of the first two iterations of coordinate descent in autotune Lasso (Algorithm 3) . The predictors are categorized into three groups: Significant predictors i.e., predictors whose true coefficients are non-zero, are depicted in green, and spurious predictors whose estimated coefficients are zero/nonzero are shown in blue and maroon bars, respectively.
In Fig.˜3, since the predictors $\text{X}_{\cdot,j}$ are uncorrelated, the SDs of the partial residuals (PRs) of significant predictors separate from the cluster of PRs of insignificant predictors right from the first iteration. We highlight in that replication, $\mathsf{autotune}$ Lasso took 30 rounds of coordinate descent to converge at the final estimate of $\beta$. In the second iteration, the separation between significant and insignificant predictors becomes wider. Now, we can carve out the support set by correctly identifying the outliers in this histogram via Algorithm 2.
We observe this phenomenon in most of the simulation setups we consider (Sec.˜4.1.1). Note that, we arrive at this by only running one round of CD on $\widehat{\beta}$ at fixed $\lambda_{0}\widehat{\sigma}^{2}$. Hence, inspired by this clustering behavior, we hypothesized that partial residuals would play a crucial role in high-dimensional regression when the sparsity assumption holds.
In general, support set recovery in Lasso is hard, especially in presence of correlated predictors [aos_support_set_recovery], where $\mathsf{autotune}$ can identify some of the insignificant predictors as outliers. But we only need PRs for $\sigma$ estimation, and even if the insignificant predictors which are strongly correlated with significant predictors, replaces the true predictors in the $\mathsf{autotune}$’s estimated support set, then also $\mathsf{autotune}$’s $\widehat{\sigma}$ doesn’t get hampered. We empirically validate this in Sec.˜4.3.1.
@@ 4 Numerical Experiments
We demonstrate the performance of $\mathsf{autotune}$ Lasso in three settings: 1) in high-dimensional regression, we compare its runtime, variable selection, estimation of $\beta$, and prediction accuracy against Scaled Lasso and CV, 2) in large VAR models, we compare its runtime, variable selection and estimation accuracy against [kock2025data] and TSCV, 3) we compare the $\widehat{\sigma}^{2}$ of autotune against Scaled Lasso, Organic and Natural Lasso, and illustrate how the PR can be used to design diagnostic plots and check the sparsity assumption.
Note that all our empirical experiments are run in Linux version 5.19.17.
@@ 4.1 High-dimensional regression
We closely follow the simulation setup of [hastie2020best], where the data-generating process is completely characterized by 6 parameters: $n$, $p$, $s$ (level of sparsity), $\rho$ (correlation between the predictors), $\mathrm{SNR}$ (Signal to Noise Ratio), and beta-type (pattern of sparsity). We use the following linear model:
@@
Following [hastie2020best], the default value for SNR is set to 2, for simulating correlated predictors, we will primarily use $\rho=0.35$ and we consider five different schemes for generating $\beta\in\mathbb{R}^{p}$:
@@
    Beta-type 3: The first $s$ entries of $\beta$ form an arithmetic progression series from 10 down to 0.5, while the rest are set to 0.
@@
    Beta-type 5: $\beta$ has $s$ nonzero entries in an arithmetic progression series from 10 down to 0.5 and placed roughly equally spaced across the index range $1$ to $p$, with all other entries set to 0.
@@
Let $\beta=\mathopen{}\mathclose{{\left(\beta_{1},\dots,\beta_{p}}}\right)$ be the true coefficient, $\widehat{\beta}=\mathopen{}\mathclose{{\left(\widehat{\beta}_{1},\dots,\widehat{\beta}_{p}}}\right)$ be the algorithm’s estimated coefficient vector, and $\text{Y}_{\mathrm{test}}|\text{X}_{\mathrm{test}}\sim\mathcal{N}\mathopen{}\mathclose{{\left(\text{X}_{\mathrm{test}}\beta,\sigma^{2}\mathbf{I}}}\right)$ where the rows of $\text{X}_{\mathrm{test}}$ are drawn from $\mathcal{N}_{p}(0,\Sigma)$, i.e., rows of X and $\text{X}_{\mathrm{test}}$ are i.i.d.. We quantify estimation and prediction error using the Relative Mean Square Error(RMSE) and Relative Test Error(RTE), respectively. RTE measures the ratio of expected prediction error using the algorithm’s estimates $\widehat{\beta}$ to that of the oracle, which knows the true $\beta$.
@@
Note that if $\widehat{\beta}=0$, i.e., the algorithm fits a null model on the data, then RMSE =1=1 and RTE $\mathrm{SNR}+1$, while for perfect fit, i.e., $\widehat{\beta}=\beta$, RMSE =0=0 and RTE =1=1. So, a good frame of reference for RMSE is $(0,1)$, while RTE varies in the range $(1,1+\mathrm{SNR})$. For variable selection, we use Area Under the Receiver Operating Characteristic curve(AUROC), a measure of how well an algortihm recovers the sparsity pattern of $\beta$.
@@
4.1.1 Effect of Autotune’s tuning in estimation accuracy
[…]
$\mathsf{autotune}$ picks the $\lambda$ closest to the RMSE-optimal $\lambda$ of Lasso solution path in both moderate and high dimensions, yielding $\widehat{\beta}$ which is almost as good as the best $\beta$ estimate one can get from the Lasso solution path after tuning via CV/AIC/BIC. On the other hand, AIC and BIC completely fail to tune the $\lambda$ in the $p>n$ regime. CV(min) and CV(1se) both overfits in moderate dimensional dataset and underfits in high dimensional dataset.
Moderate dimension, $\rho=0.35$, beta-type 1, SNR=2=2 High dimension, $\rho=0.35$, beta-type 1, SNR=2=2
Figure 4: Regularization path of autotune Lasso (blue dots) against Lasso evaluated over its default lambda grid (red dots), vertical lines indicate the final $\lambda$ selected by different procedures. Round blue dots indicate the solution path of $\mathsf{autotune}$ when $\sigma$ is being estimated (hence $\lambda=\lambda_{0}\widehat{\sigma}^{2}$ is changing) and starred blue dot display the final $\widehat{\beta}$ of $\mathsf{autotune}$.
@@ 4.1.2 Estimation, prediction and variable selection accuracy
We plot the RMSE, RTE, and AUROC of estimates $\widehat{\beta}$ given by $\mathsf{autotune}$, CV Lasso, and Scaled Lasso against a SNR grid of 20 evenly spaced values on a log scale from 0.66 to 4 in moderate and high dimensions (Tab.˜1). We report the mean of these three metrics across 200 replications along with 1 SE bars. For each setting, we considered all sparsity patterns (all beta-types) and three levels of correlation between predictors, namely $\rho\in\mathopen{}\mathclose{{\left\{0,0.35,0.7}}\right\}$ (same as in [hastie2020best]). A full suite of our simulations is provided in the Sec.˜A.2.
Our primary focus in the paper is the performance of Lasso in $n\ll p$ regime, which is realized in the high-dimensional configuration of Tab.˜1. We noticed that for SNRs below 0.66 (max PVE drops to 40%), all the methods, including autotune, are not doing any meaningful estimation (RMSE nearly 1), hence our SNR grid starts from 0.66. On the other side, we observe that all methods achieve the perfect AUROC of 1 in a large range of setups in high dimensions at SNR =4=4, so we set the maximum SNR to 4.
@@ Results:
We present a slice of the empirical comparisons in Fig.˜5 and Fig.˜6 to keep this section concise and defer the rest in Appendix (Sec.˜A.2). In this setting, $\rho=0.35$ and true coefficients follow the pattern of either beta-type 1 or beta-type 2. In beta-type 2 correlated design, the significant predictors are correlated with one another while for beta-type 1 there is practically no correlation between the significant predictors.
$\mathsf{autotune}$ uniformly shows better variable selection in terms of AUROC than CV and Scaled Lasso across the full range of setups we considered in moderate and high-dimensional settings. We also observe that in almost all settings where a benchmark method shows significant improvement in estimation or prediction accuracy over $\mathsf{autotune}$, that benchmark has an AUROC of 1 in that setup, indicating perfect model selection.
$\mathsf{autotune}$ shows significantly smaller prediction error than the benchmarks across most of the settings considered. In fact, Scaled Lasso struggles a lot in terms of RTE across all settings of high dimensional setup.
In estimation, there are 2 different scenarios. For Beta-type 1, $\mathsf{autotune}$ shows significant improvement in estimation (RMSE) over the benchmarks for all SNRs in both moderate and high-dimensional setups. Only exception was SNRs >3>3 in high dimensional setup, CV Lasso estimates better than $\mathsf{autotune}$. Scaled Lasso struggled to give any competitive performance for beta-type 1. Introducing correlation within significant predictors (Beta-type 2) reduced the RMSE of all algorithms (also observed in [reid2016study]). Scaled Lasso’s performance in estimation now matches that of $\mathsf{autotune}$ for SNRs below a critical value, after which scaled Lasso significantly becomes the best estimator. These critical values are different for moderate and high dimensions, and interestingly, scaled Lasso has AUROC of 1 for all SNRs at or beyond this critical value.
@@ 4.1.3 Runtime comparsion
We assess the speed of the algorithms through their mean runtime (in milliseconds) for estimating regression coefficients $\beta$ across 100 replications. Standard error bars across replications are also displayed in each figure.
@@
Figure 7: Runtime of $\mathsf{autotune}$, scaled Lasso and CV-Lasso with increasing $p$ for fixed $n,s$. $\mathsf{autotune}$ runs 1510015-100 times faster, with higher gain in speed in high-dimensional settings.
Table 2: Table of mean runtimes (with SD in the brackets) for Fig.˜7(a). $n=80,s=5,$ $\rho=0.35,\mathrm{SNR}=1$, Beta-type 2 and $p$ varies from 50 to 750.
@@
In Fig.˜7, we plot the mean runtimes against increasing number of predictiors $(p)$ and the values are given in Tab.˜2 and 5. $\mathsf{autotune}$ shows 15-112 times reduction in runtime as compared to CV Lasso and over 26-107 times as compared to Scaled Lasso.
4.2 Autotune VAR-Lasso
We compare $\mathsf{autotune}$ VAR against VAR Lasso tuned via TSCV, AIC, BIC, and [kock2025data] in terms of estimation, variable selection and runtime. From here on, we refer to [kock2025data] as “KPS Lasso”.
We reuse the metrics employed in the previous subsection 4.1.2, except RTE as in time series, as forecasting is interesting to econometricians and scientists while in-sample prediction is not. In variable selection, $\mathsf{autotune}$ gave AUROC of nearly 1 in all the simulation setups we considered. Hence, we use another variable selection criterion for comparison, Matthews Correlation coefficient(MCC)
$\text{MCC}=\frac{\text{TP.TN}-\text{FP.FN}}{\sqrt{(\text{TP+FP})(\text{TP+FN})(\text{TN+FN})(\text{TN+FP})}}$ (36)
where TP (and FP) is the number of true positives (and false positives) i.e., number of components of $A$ that are estimated as nonzero and are actually nonzero (zero) in the DGP. Similarly, let TN (and FN) is the number of true negatives (and false negatives). For the binary classification problem of whether $A_{i,j}=0$ or not for all $i,j$, then MCC is the correlation between the actual and estimated binary labels. So, MCC is between 1 and -1, with 1 meaning perfect classification, 0 means the classification is no better than random guessing and -1 means perfect disagreement between true and estimated labels.
@@ 4.2 Autotune VAR-Lasso
We focus on the cases $n=100$ and $200$ and we fix SNRs $\nu_{j}$ of all the $p$ time series to 2.5 for the experiments in Fig.˜8, Fig.˜9, and Fig.˜10 to keep this section concise. We depict the case of heterogenous SNRs in Fig.˜30 and Fig.˜31, where half of the $p$ time series get SNR 5, other half gets SNR 0.5, and all the other parameters of the simulation are kept same. Dimension of parameter space for both DGPs is $p^{2}$, so for $n=200$, $p=15$ gives a moderate dimensional setup and $p\geq 20$ a high dimensional setup. Our main benchmark is KPS Lasso [kock2025data], which also offers an option of post-Lasso in VAR that reduces the shrinkage bias of Lasso estimator by running “regression in two stages”. Since we are reporting the results of Lasso without any post-Lasso bias correction, to keep the comparison with $\mathsf{autotune}$ VAR fair, we use the original VAR Lasso proposed in [Algo 1, kock2025data] as KPS Lasso, instead of their post-Lasso version. We also use VAR Lasso where tuning is done via AIC, BIC and TSCV separately for each column of $\Phi$ ˜22.
@@
For estimation, $\mathsf{autotune}$’s RMSE remains fairly constant as $p$ increases, whereas RMSE of all other methods increase rapidly with dimension. This pattern becomes more prominent for smaller sample size ($n=100)$. TSCV, AIC and BIC often have RMSE >1, i.e., their estimates are worse than a null model $\widehat{A}=0_{p\times p}$. We highlight that even under such difficult high-dimensional setup, $\mathsf{autotune}$’s RMSE doesn’t explode. $\mathsf{autotune}$ uniformly dominates KPS Lasso in terms of estimation accuracy. Additionally, in $2\times 2$ block VAR DGP, KPS Lasso almost gives a null model fit for $n=100,p\geq 10$ as its RMSE becomes nearly 1, and even for $n=200$, KPS Lasso gives RMSE >0.75>0.75 for $p\geq 16$.
@@
$\mathsf{autotune}$ outperforms all the benchmarks in terms of AUROC and provides a perfect AUROC of 1 across all the setups considered in Fig.˜8 and Fig.˜9. In diagonal VAR DGP, KPS shows significantly better MCC $\mathsf{autotune}$ VAR in $n=200$ setup. In fact, KPS has the perfect MCC of 1 for $p\leq 20$. But with smaller sample size of $n=100$, MCC of KPS quickly deteriotes with increasing dimensionality of estimand $A$ and $\mathsf{autotune}$ starts significantly dominating KPS in high-dimensional regime $p^{2}>n=100$. KPS’s MCC gets significantly worse than $\mathsf{autotune}$ in $2\times 2$ block VAR DGP (Fig.˜9), it even drops below 0.2 for $p\geq 16$ in $n=100$ setup whereas $\mathsf{autotune}$ almost maintains same levels of MCC as in Diagonal VAR setup. VAR Lasso tuned by CV/AIC/BIC have much lower MCC compared to $\mathsf{autotune}$, infact their MCC drops below 0.2 for $p\geq 20$.
@@
In simulations with heterogenous SNRs (Fig.˜30, 31), the main conclusions remain same. Heterogenous SNRs further degrade the performance of VAR Lasso with TSCV, AIC, or BIC tuning in terms of RMSE and AUROC in low sample size setup of $n=100$.
@@
Figure 9: RMSE, AUROC, and Matthew Correlation Coefficient plotted as functions of number of time series $(p)$ for $n=100$ (top) and 200200 (bottom) in DGP(ii).
@@
Figure 10: Comparison of runtimes of Autotune VAR Lasso and benchmarks in two different VAR models with SNR = 2.5
In Fig.˜10, we plot the mean runtimes (with 1 SE bars) against increasing number of time series $p$ for VAR Lasso tuned with Autotune, KPS Lasso, TSCV, and AIC under the two DGPs mentioned above, BIC gave identical runtimes as AIC so we dropped it. For diagonal VAR, $\mathsf{autotune}$ is 58 - 169 times faster than TSCV and over 30 - 66 times faster than KPS Lasso. For Block VAR, $\mathsf{autotune}$ is 63 - 175 times faster TSCV and 26 - 84 times faster than KPS Lasso. AIC Lasso’s runtime was always in between KPS and TSCV VAR Lasso.
Overall, this experiment reinforces our claim that $\mathsf{autotune}$ VAR not only ameliorates running time of VAR but also improves VAR estimation in high dimensions and variable selection over KPS Lasso and popular prior works of VAR-Lasso; VAR Lasso tuned via CV, AIC, and BIC.

### v2

@@ 3.3 An illustration of PR properties:
with $n=80$ observations, $p=750$ covariates, and $\beta_{0}=(1,1,1,1,1,0,0,\dots)\in\real^{p}$, i.e., each yi depend on only the first 5 covariates. Each $\text{X}_{i,\cdot}$ is independently generated from $p$-dimensional standard normal distribution. Fixing $\text{X}_{i,\cdot}$, we generate yi by drawing errors $\epsilon_{i}$ from $\mathcal{N}(0,\sigma^{2})$ with $\sigma^{2}$ chosen so that signal-to-noise ratio is 1.
We exhibit the histogram of the SD of all the PR ˜4 obtained from the estimates of $\beta$ at the end of the first two iterations of coordinate descent in autotune Lasso (Algorithm 3) . The predictors are categorized into three groups: Significant predictors i.e., predictors whose true coefficients are non-zero, are depicted in green, and spurious predictors whose estimated coefficients are zero and nonzero are shown in blue and maroon bars respectively.
In Fig.˜3, since the predictors $\text{X}_{\cdot,j}$ are uncorrelated, the SD of the PR of significant predictors separate from the cluster of PR of insignificant predictors right from the first iteration. We highlight in that replication, $\mathsf{autotune}$ Lasso took 30 rounds of coordinate descent to converge at the final estimate of $\beta$. In the second iteration, the separation between significant and insignificant predictors becomes wider. Now, we can carve out the support set by correctly identifying the outliers in this histogram via Algorithm 2.
We observe this phenomenon in most of the simulation setups we consider. Note that, we arrive at this by only running one round of CD on $\widehat{\beta}$ at fixed $\lambda_{0}\widehat{\sigma}^{2}$. Hence, inspired by this clustering behavior, we hypothesized that partial residuals would play a crucial role in high-dimensional regression when the sparsity assumption holds.
In general, support set recovery in Lasso is known to be hard, especially in presence of correlated predictors [aos_support_set_recovery] where $\mathsf{autotune}$ can identify some of the insignificant predictors as outliers. But we only require good prediction accuracy (not support set recovery) for good $\sigma$ estimation, and even if the insignificant predictors which are strongly correlated with significant predictors are detected as outliers by $\mathsf{autotune}$, we find that $\widehat{\sigma}$ of $\mathsf{autotune}$ is reasonably accurate. We empirically validate this in Sec.˜4.3.1.
@@ 4 Numerical experiments
We demonstrate the performance of $\mathsf{autotune}$ Lasso in three settings: 1) in high-dimensional regression, we compare its runtime, variable selection, estimation of $\beta$, and prediction accuracy against Scaled Lasso and CV Lasso, 2) in large VAR models, we compare its runtime, variable selection and estimation accuracy against [kock2025data] and TSCV, 3) we compare the $\widehat{\sigma}^{2}$ of autotune against Scaled Lasso, Organic and Natural Lasso, and illustrate how the PR can be used to design diagnostic plots and check the sparsity assumption. All our empirical experiments are run in Linux version 5.19.17.
@@ 4.1 High-dimensional regression
We closely follow the simulation setup of [hastie2020best], where the data generating process (DGP) is completely characterized by 6 parameters: $n$, $p$, $s$ (level of sparsity), $\rho$ (correlation between the predictors), $\mathrm{SNR}$ (signal-to-noise ratio), and Beta-type (pattern of sparsity). We use the following linear model:
@@
Following [hastie2020best], the default value of SNR is set to 2. For simulating correlated predictors, we will primarily use $\rho=0.35$ and we consider five different schemes for generating $\beta\in\mathbb{R}^{p}$:
@@
    Beta-type 3: The first $s$ entries of $\beta$ decreases linearly from 10 to 0.5, while the rest are set to 0.
@@
    Beta-type 5: $\beta$ has $s$ nonzero entries decreasing linearly from 10 down to 0.5 and placed equally spaced across the index range $1$ to $p$, with all other entries set to 0.
@@
Let $\beta=\mathopen{}\mathclose{{\left(\beta_{1},\dots,\beta_{p}}}\right)$ be the true coefficient, $\widehat{\beta}=\mathopen{}\mathclose{{\left(\widehat{\beta}_{1},\dots,\widehat{\beta}_{p}}}\right)$ be the estimated coefficient vector of the algorithm, and $\text{Y}_{\mathrm{test}}|\text{X}_{\mathrm{test}}\sim\mathcal{N}\mathopen{}\mathclose{{\left(\text{X}_{\mathrm{test}}\beta,\sigma^{2}\mathbb{I}}}\right)$ where the rows of $\text{X}_{\mathrm{test}}$ are drawn from $\mathcal{N}_{p}(0,\Sigma)$, i.e. rows of X and $\text{X}_{\mathrm{test}}$ are i.i.d. We quantify estimation and prediction error using the Relative Mean Square Error (RMSE) and Relative Test Error (RTE), respectively. RTE measures the ratio of expected prediction error using estimates of the algorithm $\widehat{\beta}$ to that of the oracle, which knows the true $\beta$.
@@
Note that if $\widehat{\beta}=0$, i.e. the algorithm fits a null model on the data, then RMSE =1=1 and RTE $\mathrm{SNR}+1$. While for perfect fit, i.e. $\widehat{\beta}=\beta$, RMSE =0=0 and RTE =1=1. So, a good frame of reference for RMSE is $(0,1)$, while RTE varies in the range $(1,1+\mathrm{SNR})$. For variable selection, we use Area Under Receiver Operating Characteristic curve (AUROC), a measure of how well an algortihm recovers the sparsity pattern of $\beta$.
@@ 4.1.1 Estimation, prediction and variable selection accuracy
We plot the RMSE, RTE, and AUROC of estimates $\widehat{\beta}$ given by $\mathsf{autotune}$, CV Lasso, and Scaled Lasso against a SNR grid of 20 evenly spaced values on a log scale from 0.66 to 4 in moderate and high dimensions (Tab.˜1). We report the mean of these three metrics across 200 replications along with 1 standard error (SE) bars. For each setting, we considered all sparsity patterns (all Beta-types) and three levels of correlation between predictors, namely $\rho\in\mathopen{}\mathclose{{\left\{0,0.35,0.7}}\right\}$ (same as in [hastie2020best]). A full suite of our simulations is provided in the Sec.˜A.2.
Our primary focus in the paper is the performance of Lasso in the high-dimensional configuration of Tab.˜1. We noticed that for SNRs below 0.66 (max PVE drops to 40%), all the methods, including $\mathsf{autotune}$, are not doing any meaningful estimation (RMSE nearly 1), hence our SNR grid starts from 0.66. On the other side, we observe that all methods achieve the perfect AUROC of 1 in a large range of setups in high dimensions at SNR =4=4, so we set the maximum SNR to 4.
@@ Results:
We present a slice of the empirical comparisons in Fig.˜4 and Fig.˜5 to keep this section concise and defer the rest in Appendix (Sec.˜A.2). In this setting, $\rho=0.35$ and true coefficients follow the pattern of either Beta-type 1 or Beta-type 2. In Beta-type 2 correlated design, the predictors in the true support set are correlated with one another while for Beta-type 1 there is practically no correlation between the true predictors.
Our $\mathsf{autotune}$ uniformly shows better variable selection in terms of AUROC than CV and Scaled Lasso across the full range of setups we considered in moderate and high-dimensional settings. We also observe that in almost all settings where a benchmark method shows significant improvement in estimation or prediction accuracy over $\mathsf{autotune}$, that benchmark has an AUROC of 1 in that setup, indicating perfect model selection.
Our $\mathsf{autotune}$ shows significantly smaller prediction error than the benchmarks across most of the settings considered. In fact, the performance of Scaled Lasso in terms of RTE is significantly worse than $\mathsf{autotune}$ across all the setups of high-dimensional setup.
In estimation, there are 2 different scenarios. For Beta-type 1, $\mathsf{autotune}$ shows significant improvement in estimation (RMSE) over the benchmarks for all SNR in both moderate and high-dimensional setups. Only exception was SNR >3>3 in high-dimensional setup, CV Lasso estimates better than $\mathsf{autotune}$. Scaled Lasso struggled to give any competitive performance for Beta-type 1. Introducing correlation within true predictors (Beta-type 2) reduced the RMSE of all algorithms (also observed in [reid2016study]). The performance of Scaled Lasso in estimation now matches that of $\mathsf{autotune}$ for SNRs below a critical value, after which Scaled Lasso significantly becomes the best estimator. These critical values are different for moderate and high dimensions, and interestingly, Scaled Lasso has AUROC of 1 for all SNRs at or beyond this critical value.
@@
Figure 4: RMSE, RTE, and AUROC of $\mathsf{autotune}$, CV Lasso, and Scaled Lasso plotted as a function of SNR for high-dimensional setup.
@@ 4.1.2 Runtime comparsion
We assess the speed of the algorithms through their mean runtime (in milliseconds) for estimating regression coefficients $\beta$ across 100 replications. 1 SE bars across replications are also displayed in each figure.
@@
Figure 6: Runtime of $\mathsf{autotune}$, Scaled Lasso and CV Lasso with increasing $p$ for fixed $n,s$. $\mathsf{autotune}$ runs 1510015-100 times faster, with higher gain in speed in high-dimensional settings.
Table 2: Table of mean runtimes (with SD in the brackets) for Fig.˜6(a). $n=80,s=5,$ $\rho=0.35,\mathrm{SNR}=1$, Beta-type 2 and $p$ varies from 50 to 750.
@@
In Fig.˜6, we plot the mean runtimes against increasing number of predictiors $(p)$ and the values are given in Tab.˜2 and 5. $\mathsf{autotune}$ shows 15-112 times reduction in runtime as compared to CV Lasso and over 26-107 times as compared to Scaled Lasso.
4.2 VAR Lasso using $\mathsf{autotune}$
We compare VAR Lasso tuned with $\mathsf{autotune}$ against VAR Lasso tuned via TSCV, AIC, BIC, and [kock2025data] in terms of estimation, variable selection and runtime. From here on, we refer to VAR Lasso tuned via [kock2025data] as “KPS Lasso”.
Instead of out-of-sample prediction, we focus on estimation and variable selection accuracy since the primary interest in large VAR models is often learning Granger causality (lead-lag) relationships among the component time series. In variable selection, $\mathsf{autotune}$ gave AUROC of nearly 1 in all the simulation setups we considered. Hence, we use another variable selection criterion for comparison, Matthews Correlation Coefficient (MCC)
$\text{MCC}=\frac{\text{TP $\times$ TN}-\text{FP $\times$ FN}}{\sqrt{(\text{TP+FP})(\text{TP+FN})(\text{TN+FN})(\text{TN+FP})}}$ (35)
where TP (and FP) is the number of true positives (and false positives) and TN (and FN) is the number of true negatives (and false negatives). For the binary classification problem of whether $A_{i,j}=0$ or not for all $i,j$, then MCC is the correlation between the actual and estimated binary labels. So, MCC is between 1 and -1, with 1 meaning perfect classification, 0 means the classification is no better than random guessing and -1 means perfect disagreement between true and estimated labels.
@@ 4.2 VAR Lasso using $\mathsf{autotune}$
We focus on the cases $n=100$ and $200$ and we fix SNRs $\nu_{j}$ of all the $p$ time series to 2.5 for the experiments in Fig.˜7, Fig.˜8, and Fig.˜9 to keep this section concise. We depict the case of heterogenous SNRs in Fig.˜29 and Fig.˜30, where half of the $p$ time series get SNR 5, other half gets SNR 0.5, and all the other parameters of the simulation are kept same. Dimension of parameter space for both DGPs is $p^{2}$, so for $n=200$, $p=15$ gives a moderate dimensional setup and $p\geq 20$ a high-dimensional setup. Our main benchmark is KPS Lasso [kock2025data], which also offers an option of post-Lasso in VAR that reduces the shrinkage bias of Lasso estimator. Post-Lasso initially runs Lasso, collects all the predictors with non-zero $\widehat{\beta}_{j}$, and lastly reports the $\widehat{\beta}$ obtained by running OLS regression on those predictors keeping $\widehat{\beta}_{j}$ of other predictors zero. Since we are reporting the results of Lasso without any post-Lasso bias correction, to keep the comparison with $\mathsf{autotune}$ VAR fair, we use the original VAR Lasso proposed in [Algo 1, kock2025data] as KPS Lasso, instead of their post-Lasso version. We also use VAR Lasso where tuning is done via AIC, BIC and TSCV separately for each column of $\Phi$ ˜22.
@@
For estimation, RMSE of $\mathsf{autotune}$ remains fairly constant as $p$ increases, whereas RMSE of all other methods increase rapidly with dimension. This pattern becomes more prominent for smaller sample size ($n=100)$. TSCV, AIC, and BIC often have RMSE>1, i.e. their estimates are worse than a null model $\widehat{A}=0_{p\times p}$. We highlight that even under such difficult high-dimensional setup, RMSE of $\mathsf{autotune}$ does not explode. $\mathsf{autotune}$ uniformly dominates KPS Lasso in terms of estimation accuracy. Additionally, in $2\times 2$ block VAR DGP, KPS Lasso almost gives a null model fit for $n=100,$ $p\geq 10$ as its RMSE becomes nearly 1, and even for $n=200$, KPS Lasso gives RMSE >0.75>0.75 for $p\geq 16$.
@@
$\mathsf{autotune}$ outperforms all the benchmarks in terms of AUROC and provides a perfect AUROC of 1 across all the setups considered in Fig.˜7 and Fig.˜8. In diagonal VAR DGP, KPS Lasso shows significantly better MCC than $\mathsf{autotune}$ in $n=200$ setup. In fact, KPS has the perfect MCC of 1 for $p\leq 20$. But with smaller sample size of $n=100$, MCC of KPS quickly deteriotes with increasing dimensionality of estimand $A$ and $\mathsf{autotune}$ starts significantly dominating KPS in high-dimensional regime $p^{2}>n=100$. KPS’s MCC gets significantly worse than $\mathsf{autotune}$ in $2\times 2$ block VAR DGP (Fig.˜8), it even drops below 0.2 for $p\geq 16$ in $n=100$ setup whereas $\mathsf{autotune}$ almost maintains same levels of MCC as in diagonal VAR setup. VAR Lasso tuned by TSCV, AIC and BIC have much lower MCC compared to $\mathsf{autotune}$.
@@
In simulations with heterogenous SNRs (Fig.˜29, 30), the main conclusions remain same. Heterogenous SNRs further worsen the performance of VAR Lasso with TSCV, AIC, or BIC tuning in terms of RMSE and AUROC in low sample size setup of $n=100$.
@@
Figure 8: RMSE, AUROC, and Matthews Correlation Coefficient plotted as functions of number of time series $(p)$ for $n=100$ (top) and 200200 (bottom) in DGP(ii).
@@
Figure 9: Comparison of runtimes of VAR Lasso using $\mathsf{autotune}$ and benchmarks in two different VAR models with SNR = 2.5
In Fig.˜9, we plot the mean runtimes (with 1 SE bars) against increasing number of time series $p$ for VAR Lasso tuned with $\mathsf{autotune}$, KPS Lasso, TSCV, and AIC under the two DGP mentioned above, BIC gave identical runtimes as AIC so we dropped it. For diagonal VAR, $\mathsf{autotune}$ is 58 - 169 times faster than TSCV and over 30 - 66 times faster than KPS Lasso. For Block VAR, $\mathsf{autotune}$ is 63 - 175 times faster TSCV and 26 - 84 times faster than KPS Lasso. Runtime of AIC Lasso is always in between KPS and TSCV VAR Lasso.